OpenStudy (watchmath):

Compute \[\frac{3}{1\cdot 2\cdot 3}+\frac{5}{2\cdot 3\cdot 4}+\frac{7}{3\cdot 4\cdot 5}+\cdots\]

OpenStudy (anonymous):

I think the limit would converge to 1 or thereabouts.

OpenStudy (anonymous):

the general term of the expression is (2n+1)/(n(n+1)(n+2) = (n+(n+1))/(n(n+1)(n+2) then splitting the terms , =1/((n+1)(n+2)) + 1/(n(n+2)) then you can find the two individual summations to n terms then apply limit n tends to infinity.